OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of square root of 180 feet and a width of square root of 125 feet. (3 points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (jackellyn):

2(L+W)=P 2(√180+√125)=P 2(6√5+5√5)=P 2*11√5=P 22√5 ft = P

OpenStudy (anonymous):

thanks but what could the object be

OpenStudy (jackellyn):

I'm sorry, not supposed to do your homework for you :)